A mixture initially contains A, B, and C in the following concentrations: [A] = 0.700 M, [B] = 0.650 M, and [C] = 0.550 M. The following reaction occurs and equilibrium is established: A+2B<->C

At equilibrium, [A] = 0.540 M and [C] = 0.710 M. Calculate the value of the equilibrium constant, Kc.

Express your answer numerically please!
Answer #1

ICE Table:

Given at equilibrium,

[A] = 0.54

0.7-1x = 0.54

x = 0.16

Equilibrium constant expression is

Kc = [C]/[A][B]^2

Kc = (0.55+1x)/(0.7-1x)(0.65-2x)^2

Kc = (0.55+1*0.16)/(0.7-1*0.16)(0.65-2*0.16)^2

Kc = 12.1

Answer: 12.1